OpenStudy (anonymous):

What value should be added to the expression to create a perfect square? x2 – 20x

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

25, -25, 100, or -100

OpenStudy (skullpatrol):

$$\Huge x^2-20x$$

OpenStudy (anonymous):

100?

OpenStudy (skullpatrol):

yes

OpenStudy (skullpatrol):

$$\Huge x^2-20x +100=(x-10)^2$$

OpenStudy (skullpatrol):

$$\Huge a^2-2ab +b^2=(a-b)^2$$

OpenStudy (anonymous):

thanks!

OpenStudy (skullpatrol):

Thanks for asking :)